2.1 Suppose you long one Australian dollar call and one Australian dollar put with an exercise exchange rate of 0.75 (USD/AUD). The price of call and the price of put is USD0.06. Using the 15-period (spot) exchange rates given below, answer parts (a)-(c). Time Spot Exchange Rate (USD/AUD) 0 0.35 1 0.40 2 0.45 3 0.50 4 0.55 5 0.60 6 0.65 7 0.70 8 0.75 9 0.80 10 0.85 11 0.90 12 0.95 13 1.00 14 1.05 15 1.10 (a) Compute the net call payo, net put payo and net combined payo.

(b) Plot separately the Net Long call payo, the net long put payo and the net combined payo.

(c) Name and provide denition of the shape of the combined payo obtained from part (b)? 



1
Expert's answer
2021-06-09T10:23:31-0400

1a) Net call payo : ( 0.75×0.06)+(0.351×0.75)+ (0.45×1.10)

(0.045+0.26325+0.495)= 0.80325

1b) Net put payo: (0.75×0.95)× 10

0.7125×10=7.125

1c) Net combined payo: ( (0.80325+7.125)= 7.92825

B) \\\

C) half cyclic Curved